Hi, please help me solve the question:
Find the equation of a parabola passing through ( 1, -4) with the y-axis as the axis of symmetry and a vertex at (0,0).
Thank you!

Answer by lwsshak3(11628)

Find the equation of a parabola passing through ( 1, -4) with the y-axis as the axis of symmetry and a vertex at (0,0)
**
Standard form of parabola: y=A(x-h)^2+k,(h,k)=(x,y) coordinates of vertex
For given parabola:
y=Ax^2
solving for A using coordinates of given point (1,-4)
-4=A(1)^2
A=-4
Equation: y=-4x^2